OpenStudy (anonymous):

Why does t+5 squared = t + 5, but a^2 + 6a + 9 squared = |a + 3|, instead of a + 3?

OpenStudy (anonymous):

\[\sqrt{t + 5} = t + 5..\]

OpenStudy (radar):

That is the problem that you are to solve: \[\sqrt{t+5}=t+5\] Step 1. square both sides of the equation getting:\[t+5=t ^{2}+10t+25\] consolidate terms and simplify \[t ^{2}-9t+20\] quadratic equation that can be factored and solved. \[(t-5)(t-4)=0\] t=5 and t=4

OpenStudy (radar):

the other equation is stating: \[a ^{2}+6a+9=(a+3)^{2}\] which is true. Any more questions?